OpenStudy (anonymous):

Marcus has 4 3/5 pounds of cornmeal in a container. He also has a box with 1.5 pound of cornmeal in it. if he pours half the cornmeal from the box into the container, how much cornmeal will be in the container?

OpenStudy (callisto):

'' half the cornmeal from the box '' = 1/2 (1.5) = (1/2)(3/2) = 3/4 pound Originally, there are 4 3/5 pounds of cornmeal in a container So amount of cornmeal = original amount + amount added = 4 3/5 + 3/4 Can you get the answer from here?

OpenStudy (callisto):

You're welcome

OpenStudy (callisto):

I got something ugly :S 4 3/5 = 23/5 4 3/5 + 3/4 = 23/5 +3/4 = 107/20 :S